# MCTrack: A Unified 3D Multi-Object Tracking Framework for Autonomous Driving

### v1

@@ B. Implementation Details
PNAS-MOT [42] $(\textrm{arxiv}^{\prime}24)$ online 67.32 58.99 89.59 85.44 32131 2261 568 751 276
@@
Ours VirConv [67] online 80.78 84.29 89.82 86.71 32204 2188 1248 65 439
@@
RobMOT [40] $(\textrm{arxiv}^{\prime}24)$ VirConv [67] offline 81.76 85.58 91.02 86.56 32678 1714 1367 7 71
@@
Ours VirConv [67] offline 82.46 86.55 91.46 86.82 32031 2361 563 12 71
@@ 5.2 Quantitative Experiment
NuScenes: On the nuScenes dataset, MCTrack achieved an AMOTA score of 76.1%, the best performance among all participating 3D multi-object tracking systems. As shown in TABLE 2. Notably, MCTrack demonstrated superior tracking results in key detection categories such as bicycle, car, and trailer, outperforming other tracking systems. Additionally, for the Kalman filter, we employed only a simple Constant Velocity model. Moreover, MCTrack achieved the highest number of TP and the lowest number of FN and IDS. This result demonstrates MCTrack’s exceptional performance in maintaining tracking stability.
@@ 5.3 Motion Metrics Evaluation
The results show that the Kalman filter achieved the lowest VAE and VNE, indicating that it has the highest accuracy in terms of both speed magnitude and direction. Additionally, the Kalman filter also recorded the lowest VAIE, demonstrating its ability to effectively suppress speed reversal fluctuations. As expected, the differentiation method resulted in the lowest VDE, indicating the fastest speed response. The curve fitting method, on the other hand, produced the smallest VSE, meaning it generated a smoother speed curve, but at the cost of a larger VSE. As for VAIE, it is difficult to judge its quality solely based on its value, and it typically requires evaluation in the context of practical engineering applications.
@@
Detection 86721 6.50 0.59 3.78 1.49 138.71 0.07
Differentiation 86721 8.13 0.83 3.73 1.63 136.92 0.08
Curve Fitting 86721 8.64 2.48 4.33 0.73 136.83 0.10
Kalman Filter 86718 6.25 0.55 3.76 1.47 140.32 0.06
@@ A. Ro_GDIoU
MCTrack + GIoU 72.9 63.4 84911 12542 287
MCTrack + DIoU 73.0 62.9 86024 14560 282
MCTrack + Ro_GDIoU 73.7 63.8 85561 12894 331
@@
MCTrack + Ro_GDIoU 83.73 86.42 1284 1983 3
In the experiments on the nuScenes dataset, using Ro_GDIoU resulted in improvements of 0.8% and 0.7% compared to GIoU and DIoU, respectively, fully demonstrating the effectiveness of the Ro_GDIoU cost calculation strategy. However, in the experiments on the KITTI dataset, although using Ro_GDIoU also improved the HOTA metric compared to GIoU and DIoU, the improvement was not as significant as in the nuScenes dataset. We speculate that this is due to the higher detection accuracy and relatively simpler scenarios in the KITTI dataset, leading to smaller improvements in tracking performance when using Ro_GDIoU.

### v2

@@ B. Implementation Details
PNAS-MOT [41] $(\textrm{RAL}^{\prime}24)$ online 67.32 58.99 89.59 85.44 32131 2261 568 751 276
@@
Ours VirConv [66] online 80.78 84.30 89.82 86.71 32207 2185 1252 64 438
@@
Ours VirConv [66] offline 82.56 86.64 91.62 86.82 32064 2328 542 12 59
@@ 5.2 Quantitative Experiment
NuScenes: On the nuScenes dataset, MCTrack achieved an AMOTA score of 76.3%, the best performance among all participating 3D multi-object tracking systems. As shown in TABLE 2. Notably, MCTrack demonstrated superior tracking results in key detection categories such as car and trailer, outperforming other tracking systems. Additionally, for the Kalman filter, we employed only a simple Constant Velocity model. Moreover, MCTrack achieved the highest number of TP and the lowest number of FN and IDS. This result demonstrates MCTrack’s exceptional performance in maintaining tracking stability.
@@ 5.3 Motion Metrics Evaluation
The results show that the Kalman filter achieved the lowest VAE and VNE, indicating that it has the highest accuracy in terms of both speed magnitude and direction. Additionally, the Kalman filter also recorded the lowest VIR, demonstrating its ability to effectively suppress speed reversal fluctuations. As expected, the differentiation method resulted in the lowest VDE, indicating the fastest speed response. The curve fitting method, on the other hand, produced the smallest VSE, meaning it generated a smoother speed curve, but at the cost of a larger VSE. As for VAIE, it is difficult to judge its quality solely based on its value, and it typically requires evaluation in the context of practical engineering applications.
@@
Detection 88791 6.50 0.60 3.82 1.47 138.94 0.07
Differentiation 88791 8.03 0.85 3.76 1.65 137.44 0.08
Curve Fitting 88791 8.48 2.48 4.35 0.74 137.55 0.10
Kalman Filter 88692 6.30 0.56 3.78 1.46 141.04 0.06
@@ A. Ro_GDIoU
MCTrack + GIoU 73.7 63.5 84430 13851 233
MCTrack + DIoU 73.1 62.9 85231 13332 240
MCTrack + Ro_GDIoU 74.0 64.0 85900 13083 275
@@
MCTrack + Ro_GDIoU 83.90 86.42 1261 1958 3
In the experiments on the nuScenes dataset, using Ro_GDIoU resulted in improvements of 0.3% and 0.9% compared to GIoU and DIoU, respectively, fully demonstrating the effectiveness of the Ro_GDIoU cost calculation strategy. However, in the experiments on the KITTI dataset, although using Ro_GDIoU also improved the HOTA metric compared to GIoU and DIoU, the improvement was not as significant as in the nuScenes dataset. We speculate that this is due to the higher detection accuracy and relatively simpler scenarios in the KITTI dataset, leading to smaller improvements in tracking performance when using Ro_GDIoU.